# Sequential Quadratic Sum-of-squares Programming for Nonlinear Control Systems

### v1

@@ Abstract
Many problems in nonlinear systems analysis and control design, such as local region-of-attraction estimation, inner-approximations of reachable sets or control design under state and control constraints can be formulated as nonconvex sum-of-squares programs. Yet tractable and efficient solution methods are still lacking, limiting their application in control engineering. To address this gap, we propose a filter line-search algorithm that solves a sequence of quadratic subproblems. Numerical benchmarks demonstrate that the algorithm can significantly reduce the number of iterations, resulting in a substantial decrease in computation time compared to established methods for nonconvex sum-of-squares programs. An open-source implementation of the algorithm along with the numerical benchmarks is provided.
Index Terms:
Constraint control, Lyapunov methods, Numerical algorithms, Optimization algorithms, Sum-of-squares optimization, Variational methods
I Introduction
Sum-of-squares (SOS) optimization provides a computationally tractable framework for certifying polynomial nonnegativity [42]. For nonlinear systems with polynomial dynamics, this property is powerful as many problems arising in nonlinear system analysis and control design can be reformulated into the question of nonnegativity by the Positivstellensatz [47] and thus solved by SOS programming. For convex problems, for example, certification of global asymptotic stability of nonlinear systems, the SOS problem can be transcribed into and solved by semi-definite programs (SDPs). Depending on the problem size, this powerful tool comes with potential high-computational effort. Over the past decades substantial improvement in computational performance can be observed for convex SOS problems, not only due to the improvements in computing hardware. This includes efficient conic solvers such as [4, 22, 38, 21, 23], specialized SOS solver [41], making use of relaxations to (scaled) diagonally dominant cones [2], anytime algorithms [17], efficient toolboxes in different programming languages, among others, [40, 29, 14] or, exploiting the problem structure, such as symmetry [31] or sparsity [1].
@@ I Introduction
However, many practical control engineering problems—such as stability, invariance, and controllability—typically hold only over local regions of the state space [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]. Verifying these properties within such local regions leads to nonconvex SOS optimization problems, which prohibits the direct usage of SDPs. Compared to the convex case, less improvement has been made for the nonconvex cases. State-of-the art methods to deal with the non-convexity are iterative schemes [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]. These schemes include bisection for quasiconvex problems (see [46] for details), coordinate descent, sometimes also referred as alternating-direction, with convex subproblems, and hybrid approaches combining coordinate descent with bisections for quasiconvex subproblems. Table I shows which method is used for different control applications. Coordinate descent iteratively optimizes over one variable at a time while keeping others fixed, simplifying nonconvex problems into sequences of convex SOS programs. However, unless the underlying problem is quasiconvex, these techniques lack convergence guarantees, may require many iterations with unpredictable performance and require a feasible initial guess. Additionally, more effort is needed by the user to split the nonconvex problem into the convex subproblems. There also exist nonlinear SDP solver, e.g., [19], yet their development seems to be stalled and hence is not further considered.
@@
TABLE I: Overview for different control applications and their solution method using either coordinate-descent for convex subproblems (C-CD) or a hybrid approaches (h-CD) to deal with quasiconvex subproblems.
Application Ref. H-CD C-CD
Region-of-attraction [8, 9, 24, 32]
Reachability [52, 53, 26]
[11, 33, 30]
Control-synthesis [26, 34]
Control Barrier Function [3, 10]
Local Control Lyapunov Function [48]
Control Barrier Function & [45]
Control Lyapunov Function
Nonlinear Observer [48]
Inspired by sequential convex programming, we proposed sequential (linear) SOS programming [12]. This method is a promising candidate to reduce the number of iterations compared to coordinate descent and provides (local) convergence guarantees. However, it lacks a comprehensive globalization strategy, does not explicitly address constraint violations, and it lacks an available efficient implementation for practical usage. Furthermore, for practical problems faster convergence is desirable. Thus, leveraging a sequential quadratic programming approach with a suitable globalization strategy is desirable.
Efficient constraint violation checks are crucial, particularly in nonlinear optimization. While simple function evaluations in combination with some norm are used in parameter optimization, conic and SOS programming require projections onto the corresponding cone(s) [7]. As no analytical projection operator is available for the SOS cone, projections must be computed using reliable numerical methods
Globalization strategies are critical for ensuring global convergence in Newton’s method [5, Ch.1]. Merit functions, such as $l_{p}$-norm penalty or augmented Lagrangian, balance cost reduction and constraint violation but require careful penalty parameter selection. This can introduce additional computational costs due to the need for iterative penalty adjustments (cf. [36, 6]) and introduce numerical instability if wrongly selected. This issue is especially relevant in the context of SOS constraint violations, as frequent checks are needed for the penalty adjustment, which may significantly increase the computational burden. Trust-region methods offer superior convergence properties [6], yet, they come at a higher computational cost and lead to larger problems due to additional (norm) constraints.
The filter algorithm [20] offers a promising alternative by decoupling feasibility and optimality. It tracks objective values and constraint violations, making it robust to penalty parameter sensitivity and gives more flexibility in selecting the next solution candidate [6]. Unlike Merit functions, the filter explicitly handles infeasibility through a restoration phase [20, Section 3.3], alleviating numerical instability and ensuring global convergence at the cost of a higher implementation effort.
@@
Given these considerations, a filter line-search algorithm appears to be a suitable candidate for the SOS context. We therefore propose such an approach, which solves a sequence of quadratic subproblems and thereby significantly enhances the method introduced in [12]. Our work builds upon an established filter line-search algorithm proposed in [50] and also takes inspiration from [51]. Whereas global convergence for an active set method is proven in [50], this cannot be extended to SOS in a straightforward manner. We leave the global convergence proof for future work and concentrate on the practical aspects and provide a local convergence analysis.
The key contributions of this paper are as follows.
Convergence Analysis
@@
A local convergence analysis for the sequential SOS algorithm with quadratic subproblems is provided.
Design Aspects
@@
SOS specific design aspects for nonlinear programming are investigated and discussed including the efficient constraint violation check and the design of a feasibility restoration phase.
Practical Implementation
@@
An open-source implementation of the proposed filter line-search algorithm is provided.
@@
In several case studies for real-world inspired examples, we demonstrate the effectiveness and practical applicability of the proposed approach for various control problems. Additionally, we compare it to state-of-the-art methods for nonconvex SOS programming.
II Problem Statement
Motivational Example
We start with a practical problem arising in control engineering. Assume a continuous-time nonlinear dynamic system defined by a polynomial vector field $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}[x]^{n}$, where $x\in\mathbb{R}^{n}$ denotes the state vector and $u\in\mathbb{R}^{m}$ is the control input, satisfying $f(0,0)=0$. We are interested to synthesize a (linear) control law $\kappa(x)$ which asymptotically stabilizes the system’s origin in a domain $\mathcal{X}=\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$, i.e., we are interested in a non-conservative state-constrained region-of-attraction. We therefore want to maximize the volume of the sublevel set of the ROA while fulfilling two set-containment constraints, i.e.,
$\max\quad$ $\mathrm{vol}(V,\gamma)$
s.t. $\{x\in\mathbb{R}^{n}\mid V\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$
$\{x\in\mathbb{R}^{n}\mid V\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid\langle\nabla V,f_{\kappa}\rangle<0\}$
where $\mathrm{vol}(V,\gamma)$ is the Lebesque measure of the $\gamma$-sublevel set of $V$, $f_{\kappa}=f(\cdot,\kappa(\cdot))$ are the closed-loop polynomial dynamics, $\langle\cdot,\cdot\rangle$ denotes the inner product, $V>0$ is the sought Lyapunov function, and $\gamma$ is a stable level of $V$. Statements of forward invariance, asymptotic stability, regions of attraction, passivity, bounds on the $L_{2}$-gain, input-to-state stability, control Lyapunov and barrier functions, and many more properties of nonlinear control systems can be written as dissipation inequalties [18]. Sum-of-squares constraints provide systematic formulations of dissipation inequalities for polynomial dynamical systems.
II-A Sum-of-squares Polynomials
@@
Let $x=(x_{1},\ldots,x_{n})$ be a tuple of indeterminate variables and $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}_{0}^{n}$ a multi-index. A polynomial $p$ in $x$ up to degree $d$ is a linear combination
$p=\sum_{|\alpha|_{1}\leq d}c_{\alpha}x^{\alpha}$
@@
where $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ and $\|\alpha\|_{1}=\sum_{i}\alpha_{i}$. Denote $\mathbb{R}[x]$ as the set of polynomials in $x$ with real coefficients $c_{\alpha}\in\mathbb{R}$ up to degree $d$.
Definition 1.
@@
The cone of sum-of-squares $\Sigma_{d}[x]$ forms a convex cone in $\mathbb{R}_{d}[x]$ with its dual cone $\Sigma_{d}[x]^{*}$, which is isometric to the cone of sum-of-squares polynomials [28].
As stated in [42, Theorem 3.3] a convex sum-of-squares optimization problem can be reduced to an SDP. Given $p\in\mathbb{R}_{d}[x]$, $s_{d}\in\mathbb{N}$ denotes the number of monomials up to degree $d$ of $p$. A polynomial $p\in\mathbb{R}_{2d}[x]$ is sum-of-squares if and only if there exists a matrix $Q\in\mathbb{S}_{s_{d}}^{+}$ satisfying $p=\zeta^{\top}Q\zeta$, where $\zeta\in\mathbb{R}_{d}[x]^{s_{d}}$ is the vector of monomials up to degree $d$, and $\mathbb{S}_{s_{d}}^{+}$ denotes the set of positive semidefinite matrices.
To transcribe the high-level optimization problem into a SOS problem we make use of the so-called generalized $\mathcal{S}$-procedure, a relaxed version of the Positivstellensatz [48].
Theorem 1 ([48]).
holds.
II-B Nonconvex Sum-of-squares Optimization
Applying Theorem 1 to the motivational example yields
s.t. $s_{1}(V-\gamma)-g(x)$ $\in\Sigma[x]$
$V-\epsilon(x)$ $\in\Sigma[x]$
where $s_{1},s_{2}$ are SOS multipliers resulting from Theorem 1. One can observe that the problem is bilinear in the decision variables $s_{1},s_{2},V$ and $\kappa$ is entering nonlinearly into the second constraint. Clearly, this problem cannot be directly transcribed into and solved by an SDP.
We are interested to efficiently solve general nonconvex, possibly nonlinear SOS problems of the form
@@ II-B Nonconvex Sum-of-squares Optimization
where $f:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $g:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are differentiable functions. Since the degrees $d,d^{\prime}\in\mathbb{N}_{0}$ are fixed, $f$ and $g$ are mappings between finite vector spaces and the gradients $\nabla f(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $\nabla g(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are finite linear operators. Define the Lagrangian as $L(\xi,\lambda)=f(\xi)-\langle\lambda,g(\xi)\rangle$} with Lagrange variable $\lambda\in(\Sigma_{2d^{\prime}}[x]^{m})^{*}$. Under a suitable constraint qualification, the optimal solution $\xi_{\star}$ of (1) satisfies the Karush–Kuhn–Tucker (KKT) necessary conditions
@@
III Methodology
This section provides the general concept and core idea of the proposed approach. We start with the basic principle behind sequential quadratic (SOS) programming. Unlike [12], we consider quadratic instead of linear subproblems. A local convergence analysis for quadratic problems is provided. Finally, the concept of filter line-search as globalization strategy is introduced which serves as basis for the practical implementation.
III-A Sequential Quadratic SOS Programming
We solve the nonlinear optimization problem (1) by a sequence of local convex sum-of-squares problems relative to a previous candidate solution (or initial guess) $\xi^{k}$. The solution of the local problem is then used to update the solution candidate and another local problem is solved, as shown in Alg. 1. This framework is a standard approach in nonlinear optimization [36].

### v2

@@ Abstract
tractable framework for certifying polynomial nonnegativity. If the considered problem is convex, the SOS problem can be transcribed into and solved by semi-definite programs. However, in case of nonconvex problems iterative procedures are needed. Yet tractable and efficient solution methods are still lacking, limiting their application, for instance, in control engineering. To address this gap, we propose a filter line search algorithm that solves a sequence of quadratic subproblems. Numerical benchmarks demonstrate that the algorithm can significantly reduce the number of iterations, resulting in a substantial decrease in computation time compared to established methods for nonconvex SOS programs. An open-source implementation of the algorithm along with the numerical benchmarks is provided.
journal: Mathematical Programming Computation
1 Introduction
Checking global nonnegativity of a function of several variables is a fundamental question that arises in many areas of applied mathematics. A tractable way to answer this question lies in polynomials that can be written as a sum of squares, a strict subset of the nonnegative polynomials. In parillo2003 it was shown that convex optimization over sum-of-squares (SOS) polynomials can be transcribed into and solved by semi-definite programs (SDPs). SOS programming problems arise from polynomial optimization blekherman2013 , robust and stochastic optimization bertsimas2011 , statistics and machine learning magnani2005 ; barak2015 , and control engineering jarvis-wloszek2003 ; jarvis-wloszek2005 , to name a few domains and applications.
Depending on the problem size, this powerful tool comes with potentially high computational efforts. Over the past decades substantial improvement in computational performance can be observed for convex SOS problems, not only due to the advances in computing hardware. This includes improved conic solvers mosek ; Clarabel_2024 ; scs ; operations ; han2024 , specialized SOS solvers papp2022 , use of relaxations to (scaled) diagonally dominant cones ahmadiMajumdar2017 , anytime feasible algorithms anysos , efficient implementations of polynomial operations and the reduction of a SOS program to an SDP jagt_efficient_2022 ; legat2017 ; Cunis2025 , or exploiting symmetry loefberg2009 and sparsity ahmadiEtAl2017 in the problem structure.
1.1 Nonconvex Sum-of-Squares Problems
For nonconvex SOS problems less advancement can be observed. In the context of systems and control theory, many practical engineering problems – such as stability, invariance, and controllability – typically hold only over local regions of the state space chakrabortyEtAl2011 ; chakrabortyEtAl2011ii ; iannelli2018 ; seilerBalas2010 ; yinEtAl2019 ; Yin2021 ; jarvisEtAl2003 ; Cunis2021aut ; majumdarEtAl2013 ; lin2022 ; newton2025a ; amesControlBarrierFunctions2019 ; clark2021 ; tan2006 ; schneeberger_sos_2023 ; loureiro2025 . Verifying these properties within such local regions leads to nonconvex SOS optimization problems, which prohibits the direct usage of (convex) SDPs. While nonconvex problem might also arise in other domains, they seem to have attracted little interest, possibly due to the hitherto lack of suitable solvers.
@@ 1.1 Nonconvex Sum-of-Squares Problems
State-of-the art methods for nonconvex problems are iterative schemes chakrabortyEtAl2011 ; chakrabortyEtAl2011ii ; iannelli2018 ; seilerBalas2010 ; yinEtAl2019 ; Yin2021 ; jarvisEtAl2003 ; Cunis2021aut ; majumdarEtAl2013 ; lin2022 ; newton2025a ; amesControlBarrierFunctions2019 ; clark2021 ; tan2006 ; schneeberger_sos_2023 ; loureiro2025 . These schemes include bisection for quasiconvex problems (see seilerBalas2010 for details), coordinate descent (also referred to as alternating-direction) with convex subproblems, and hybrid approaches combining coordinate descent with bisections for quasiconvex subproblems. Table 1 lists different control applications and the iterative methods applied. Coordinate descent iteratively optimizes over one variable at a time while keeping others fixed, simplifying nonconvex problems into sequences of convex SOS programs. However, unless the problem is quasiconvex, these techniques lack convergence guarantees, may require many iterations with unpredictable performance, and require a feasible initial guess. Additionally, more effort is needed by the user to split the nonconvex problem into the convex subproblems. There also exist nonlinear SDP solvers such as PENLAB fiala2013 , yet their development appears to have stalled and hence is not further considered.
@@
Table 1: Overview of control applications with their solution method using coordinate-descent with convex subproblems or a hybrid approach with quasiconvex subproblems.
Application Ref. hybrid convex
Region-of-attraction analysis chakrabortyEtAl2011 ; chakrabortyEtAl2011ii ; iannelli2018 ; loureiro2025
Reachability analysis yinEtAl2019 ; Yin2021 ; jarvisEtAl2003
Cunis2021aut ; majumdarEtAl2013 ; lin2022
Feedback law synthesis jarvisEtAl2003 ; newton2025a
Control barrier function synthesis amesControlBarrierFunctions2019 ; clark2021
Local control Lyapunov function synthesis tan2006
Control barrier & control Lyapunov function synthesis schneeberger_sos_2023
Nonlinear observer design tan2006
1.2 Challenges for a Practical Implementation
Inspired by sequential convex programming, we proposed sequential (linear) SOS programming Cunis2023acc . This method reduced the number of iterations compared to coordinate descent and provides (local, linear) convergence guarantees as an instance of quasi-Newton methods. However, it lacks a comprehensive globalization strategy and it does not explicitly address constraint violations. Furthermore, for practical problems faster convergence is desirable. Thus, leveraging a sequential quadratic, instead of sequential linear programming approach with a suitable globalization strategy is desirable. Efficient constraint violation checks are crucial, particularly in nonlinear optimization used for system analysis and verification, to ensure correctness of the results. While simple function evaluations in combination with some norm are used in parameter optimization, conic and SOS programming require projections onto the corresponding cone(s) Boyd2005 . As no analytical projection operator is available for the SOS cone, projections must be computed using reliable numerical methods. Globalization strategies are critical for ensuring global convergence of Newton-type methods (betts2010, , Chapter 1). Merit functions, such as $l_{p}$-norm penalty or augmented Lagrangian, in combination with line search, balance cost reduction and constraint violation but require careful penalty parameter selection. This can introduce additional computational costs due to the need for iterative penalty adjustments (cf. nocedalNumericalOptimization2006 ; bieglerNonlinearProgrammingConcepts2010 ) and introduce numerical instability if wrongly selected. This issue is especially relevant in the context of SOS constraint violations as frequent checks, needed for the penalty adjustment, may significantly increase the computational burden. The filter algorithm fletcherNonlinearProgrammingPenalty2002 offers a promising alternative by decoupling feasibility and optimality. It tracks objective values and constraint violations, making it robust to penalty parameter selection and gives more flexibility in computing the next solution candidate bieglerNonlinearProgrammingConcepts2010 . Unlike Merit functions, the filter explicitly handles infeasibility through a restoration phase (fletcherNonlinearProgrammingPenalty2002, , Section 3.3), alleviating numerical instability and ensuring global convergence at the cost of a higher implementation effort. Unlike line search, trust-region methods modify both the step-length and search direction using either Merit functions or a filter. According to (bieglerNonlinearProgrammingConcepts2010, , Section 5.7) they offer superior convergence properties (compared to line search), yet, they come at a higher computational cost and lead to larger problems due to additional (norm) constraints. Given these considerations, a filter line search algorithm appears to be a suitable candidate for nonconvex SOS programs.
1.3 Contribution and Structure
@@
Based on the considerations in the previous sections, we propose a filter line search approach to sequential quadratic SOS programming which significantly enhances the method introduced in Cunis2023acc . Our work builds upon the filter line search algorithm established in wachterLineSearchFilter2005 while takking inspiration from wachterImplementationInteriorpointFilter2006a . Whereas global convergence for an active set method is proven in wachterLineSearchFilter2005 , the extensions to SOS is not straightforward. We leave the global convergence analysis for future work, concentrating on the practical aspects and providing a local convergence proof. The key contributions of this paper are as follows:
@@
  • Convergence Analysis: A local convergence analysis for the sequential SOS algorithm with quadratic subproblems is provided.
@@
  • Design Aspects: SOS-specific design aspects for nonlinear programming are investigated and discussed including the efficient constraint violation check and the design of a feasibility restoration phase.
@@
  • Practical Implementation: An implementation of the proposed filter line search algorithm is provided as part of the open-source software CaΣoS Cunis2025a .
@@
We demonstrate the effectiveness and practical applicability of the proposed approach in several benchmark problems inspired from real-world control engineering examples. Additionally, we compare it to state-of-the-art methods for nonconvex SOS programming.
This paper is structured as follows: In Section 2 the problem statement of nonconvex SOS problems is motivated and explained. A proof for local convergence and the background on the filter line search algorithm are provided in Section 3. The details regarding the practical implementation are given in Section 4. In Section 5 the proposed algorithm is benchmarked against the state-of-the art solution method for nonconvex SOS programs. Practical aspects are discussed at the end.
2 Problem Statement
2.1 Sum-of-squares Polynomials
@@
Let $x=(x_{1},\ldots,x_{n})$ be a tuple of indeterminate variables. A polynomial $p$ in $x$ up to degree $d$ is a linear combination
$p=\sum_{\|\alpha\|_{1}\leq d}c_{\alpha}x^{\alpha}$
@@
with multi-indices $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}_{0}^{n}$, where $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ and $\|\alpha\|_{1}=\sum_{i}\alpha_{i}$. Denote by $\mathbb{R}_{d}[x]$ the set of polynomials in $x$ with real coefficients $c_{\alpha}\in\mathbb{R}$ up to degree $d$.
Definition 1
@@
The cone of sum-of-squares polynomials $\Sigma_{d}[x]$ forms a convex cone in $\mathbb{R}_{d}[x]$ with dual cone $\Sigma_{d}[x]^{*}$, which is isometric to the cone of sum-of-squares polynomials Lasserre2001 . Given $d\in\mathbb{N}$, denote by $s_{d}\in\mathbb{N}$ the number of monomials in $x$ up to degree $d$. A polynomial $p\in\mathbb{R}_{2d}[x]$ is sum-of-squares if and only if there exists a positive semidefinite matrix $Q\in\mathbb{R}^{s_{d}\times s_{d}}$ satisfying $p=\zeta^{\top}Q\zeta$, where $\zeta\in\mathbb{R}_{d}[x]^{s_{d}}$ is the vector of monomials up to degree $d$.
2.2 Nonconvex Sum-of-squares Optimization
We are interested in solving general nonconvex SOS problems of the form
@@ 2.2 Nonconvex Sum-of-squares Optimization
where $f:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $g:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are (twice) differentiable functions. Since the degrees $d,d^{\prime}\in\mathbb{N}_{0}$ are fixed, $f$ and $g$ are mappings between finite vector spaces and the gradients $\nabla f(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $\nabla g(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are finite linear operators. Define the Lagrangian as $L(\xi,\lambda)=f(\xi)-\langle\lambda,g(\xi)\rangle$} with Lagrange variable $\lambda\in(\Sigma_{2d^{\prime}}[x]^{m})^{*}$. Under a suitable constraint qualification, the optimal solution $\xi_{\star}$ of (1) satisfies the Karush–Kuhn–Tucker (KKT) necessary conditions
@@
2.3 Motivational Example
We give a practical problem arising in control engineering. Assume a continuous-time nonlinear dynamic system defined by a polynomial vector field $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}[x]^{n}$, where $x\in\mathbb{R}^{n}$ denotes the state vector and $u\in\mathbb{R}^{m}$ is the control input, satisfying $f(0,0)=0$. We are interested to synthesize a (linear) control law $\kappa(x)$ which asymptotically stabilizes the system’s origin in a domain $\mathcal{X}=\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$, i.e., we are interested in a non-conservative state-constrained region-of-attraction. We therefore want to maximize the volume of the sublevel set of the ROA while fulfilling two set-containment constraints, i.e.,
$\max_{V,\kappa}\quad$ $\mathrm{vol}(V,\gamma)$ (3a)
s.t. $\{x\in\mathbb{R}^{n}\mid V(x)\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$ (3b)
$\{x\in\mathbb{R}^{n}\mid V(x)\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid\langle\nabla V(x),f_{\kappa}(x)\rangle<0\}$ (3c)
where $\mathrm{vol}(V,\gamma)$ is the Lebesque measure of the $\gamma$-sublevel set of $V$, $f_{\kappa}=f(\cdot,\kappa(\cdot))$ are the closed-loop polynomial dynamics, $\langle\cdot,\cdot\rangle$ denotes the inner product, $V>0$ is the Lyapunov function, and $\gamma$ is a stable level of $V$.
To transcribe the high-level problem (3) into the SOS problem (1), we make use of the so-called generalized S-procedure, a special case of the Positivstellensatz.
Theorem 2.1(Generalized S-procedure tan2006 )
holds.
Applying Theorem 2.1 to the motivational example (3) yields
s.t. $s_{1}(V-\gamma)-g(x)$ $\in\Sigma[x]$
$V-\epsilon(x)$ $\in\Sigma[x]$
where $s_{1},s_{2}$ are SOS multipliers resulting from Theorem 2.1 and $\epsilon(x)$ is a small term to ensure stric inequalities are met. One can observe that the problem is bilinear in the decision variables $s_{1},s_{2},V$ and $\kappa$ is entering nonlinearly into the second constraint. Clearly, this problem cannot be directly transcribed into and solved by an SDP.
Statements of forward invariance, asymptotic stability, regions of attraction, passivity, bounds on the $L_{2}$-gain, input-to-state stability, control Lyapunov and barrier functions, and many more properties of nonlinear control systems can be written as dissipation inequalities ebenbauerAllgoewer2006 of the form
$\nabla\phi(x)f(x,u)\leq\varsigma(x,u)\quad\forall(x,u)\in\Omega\times\mathcal{U}$
where $\phi$ and $\varsigma$ are the so-called storage function and supply rate, $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ are the system dynamics, and $\Omega\subset\mathbb{R}^{n}$ and $\mathcal{U}\subset\mathbb{R}^{m}$ are subsets of the state and input space, respectively. Sum-of-squares constraints provide systematic formulations of dissipation inequalities for polynomial dynamical systems and play an important role for the benchmarks in Section 5.
3 Methodology
This section provides the general concept and core idea of the proposed approach. We start with the basic principle behind sequential quadratic (SOS) programming. Unlike Cunis2023acc , we consider quadratic instead of linear subproblems. A local convergence analysis for quadratic problems is provided. Finally, the concept of filter line search as globalization strategy is introduced which serves as basis for the practical implementation.
3.1 Sequential Quadratic SOS Programming
We solve the nonlinear optimization problem (1) by a sequence of local convex sum-of-squares problems relative to a previous candidate solution (or initial guess) $\xi^{k}$. The solution of the local problem is then used to update the solution candidate and another local problem is solved, as shown in Alg. 1. This framework is a standard approach in nonlinear optimization nocedalNumericalOptimization2006 .